Need some help here. 0U nd 5 sinks. 12. Refer to Table: Production. Which of the following statements is correct? The carpenter has an absolute an
Answer #1

s pz 100 -1009 ps 10+5gs At equilibrium po PS 100 - 1080 - 10+593 At equilibrium go-ge 100-109 = 10 +59 90 = 159 8 = 6 unill